OpenStudy (anonymous):

solve this equation for y 5x - 7y = 10 a. y= -7/5x - 7/10 b. y= 7/5x + 10/5 c. y- 5/7x - 10/7 d. y= -5x - 7

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

5x - 7y = 10 5x-10=7y y=?

OpenStudy (anonymous):

@theopenstudyowl

OpenStudy (nurali):

Divide by 7

OpenStudy (anonymous):

You typed faster than me :( XD @Nurali

OpenStudy (anonymous):

Im still lost.

OpenStudy (anonymous):

You got it to \[5x-10=7y\] You have to divide both sides by 7, What you do to one side, you have to do to the other side.

OpenStudy (anonymous):

Then you must isolate Y So \[5x-7y-10\]

OpenStudy (anonymous):

\[-7y=-5x-10\] You have to change \[5x\] to this side as negative.

OpenStudy (anonymous):

Then divide by -7 and you end up with \[y= 5/7x-10/7\] So theres your answer! Any questions?